Karen works for $10 an hour. A total of 25% of her salary is deducted for taxes and insurance. She is trying to save $450 for a new set of tires. Write an equation to help determine how many hours she must work to take home $450 if she saves all of her earnings. Be sure to save the equation you write, as you will use it to answer the next question in this quiz.


10h - .25=450

10h + .25(10h)=450

h -.25 (10h)=450

10h - .25(10h)=450


Answer by addingup(3677)

h{10-(0.25*10}
= h(10-2.50)
= h(7.50) = her take-home pay Now solve this:
h(7.5) = 450
h = 450/7.5 divide to get how many hours she has to work for the 450